OpenStudy (anonymous):

Would somebody please show me how to solve this type of problem? Midpoint of line segments.

OpenStudy (anonymous):

Here is the one I'm working on

OpenStudy (anonymous):

I know the formula but I'm not sure of how to apply it I believe it was (x1+x2)/2, (y1+y2)/2?

OpenStudy (jdoe0001):

\(\large \textit{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\\ \quad \\ \quad \\ \begin{array}{lllllll} &&x_1&y_1&x_2&y_2\\ &T(&2,&2)\qquad\qquad U(&11,&-7) \end{array} \\ \quad \\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\implies \left(\cfrac{(11) + (2)}{2}\quad ,\quad \cfrac{(-7) + (2)}{2} \right)\)

OpenStudy (anonymous):

Oh okay I see how i'm supposed to set it up now, much easier with an example. Thanks so much!

OpenStudy (jdoe0001):

yw